• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 08:03:33 on February22,2002

    Entry number 3513

    keyword=Owl shift summary

    02:10 Temporarily secured RCS DAQ troubleshooting efforts and restored hall to Beam Permit. Ting will resume with the help of Bob Michaels in a few hours. We are making preparations for a Moller measurement and e-p measurement.
    03:43 Completed Moller data-taking; result 78.13%
    05:30 Completed ep data-taking. Bodo will make a HALOG entry with more details, but it seems he saw both a potential problem with detector voltages and possibly the necessity to replace the ep target. See his entry for details. We couldn't get hold of the ARC measurement guys so we will just do optics runs for a while.
    05:45 Started taking optics runs at 20 uA.
    06:08 Had to reset RCS HV crate 1.
    06:10 Beam taken away for daily q.e. measurement.
    06:20 Beam is back. Continuing optics runs.
    08:80 Optics, optics, optics...